if the numerator of a fraction is decreased by 1,its value become 1/3.if the denominator of it is increased by 3,the value becomes 1/3.what is the fraction?
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

if the numerator of a fraction is decreased by 1,its value become 1/3.if the denominator of it is increased by 3,the value becomes 1/3.what is the fraction?
----
Let the fraction be a/b
-----
Equations:
(a-1)/b = 1/3
a/(b+3) = 1/3
-------------------------
Rearrange:
3a -3 = b
3a = b + 3
------------------
Rearrange for elimination:
3a - b = 3
3a - b = 3
-------------------
b = 3a-3
-------------------
Answer:: The fraction is a/(3(a-1)) where a is any number except 1.
----------------
Cheers,
Stan H.
---------------

Answer by MathTherapy(10552)

if the numerator of a fraction is decreased by 1,its value become 1/3.if the denominator of it is increased by 3,the value becomes 1/3.what is the fraction?
The fraction is any numerator-integer (N) except 1, and a denominator that is equal to 3N – 3.